# BEAST: Efficient Tokenization of B-Splines Encoded Action Sequences for Imitation Learning

### v2

@@ 6 Conclusion
We present BEAST, a B-spline–based tokenizer for continuous robot actions that compresses arbitrary trajectories into fixed-length token sequences while preserving smooth transitions between segments. BEAST supports discrete and continuous outputs and integrates seamlessly with various model architectures. By exploiting parallel decoding, it delivers fast inference and high compression rates without sacrificing performance. In extensive experiments—both in simulation and on real robots—BEAST consistently achieves strong results, demonstrating the effectiveness of our tokenization strategy. Limitations: Although BEAST delivers strong performance, it is sensitive to the choice of the number of B-spline basis functions, which can markedly affect task outcomes (Section 5.5). The optimal count depends on the smoothness and sampling frequency of the trajectory; our experiments indicate that using 5–10 bases works well for one-second robot trajectories. Future Work: We plan to extend BEAST to large-scale robot pretraining and to integrate continuous token representations with diffusion- and flow-matching objectives, aiming to further boost downstream task performance.
Appendix A Architectures
@@
Appendix B Baselines Implementation
@@
Appendix C Hyperparameters
@@
Appendix D Real Robots Setup & Tasks
D.1 Robot System Details
@@
D.2 Tasks Description and Evaluation Metrics
@@
Appendix E Compute Resources

### v3

@@ 6 Conclusion
We present BEAST, a B-spline–based tokenizer for continuous robot actions that compresses arbitrary trajectories into fixed-length token sequences while preserving smooth transitions between segments. BEAST supports discrete and continuous outputs and integrates seamlessly with various model architectures. By exploiting parallel decoding, it delivers fast inference and high compression rates without sacrificing performance. In extensive experiments—both in simulation and on real robots—BEAST consistently achieves strong results, demonstrating the effectiveness of our tokenization strategy.
Limitations: Although BEAST delivers strong performance, it is sensitive to the choice of the number of B-spline basis functions, which can markedly affect task outcomes (Section 5.6). The optimal count depends on the smoothness and sampling frequency of the trajectory; our experiments indicate that using 5–10 bases works well for one-second robot trajectories. A heuristic for selecting the number of basis functions is provided in Appendix A.
Future Work: We plan to extend BEAST to large-scale robot pretraining and to integrate continuous token representations with diffusion- and flow-matching objectives, aiming to further boost downstream task performance. Preliminary results on combining continuous BEAST tokens with a flow-matching objective are provided in Appendix D.
7 Acknowledgment
This work was supported by the pilot program Core Informatics of the Helmholtz Association (HGF), by the German Research Foundation (DFG, grant no. 448648559), and in part by the Helmholtz Association of German Research Centers. The authors also acknowledge support by the state of Baden-Württemberg through the HoreKa supercomputer funded by the Ministry of Science, Research and the Arts Baden-Württemberg, and by the German Federal Ministry of Education and Research.
Appendix A Choice of Number of Basis Functions
For finding the optimal N, we recommend starting with a number of basis functions equal to half of the sequence length. In practice, we typically sample a batch of action sequences from the dataset (e.g., 100 trajectories or more) and compute the reconstruction mean square error (MSE). An MSE below 1e-2 (for action sequences that are normalized into the range [1,1][-1,1]) usually indicates that the chosen number of basis functions allows BEAST to represent the original trajectories well. We then repeat this process to find the minimal number of basis functions that achieves this threshold, thereby maximizing compression. Since BEAST does not require separate tokenizer training, this hyperparameter tuning process can be done in a few minutes across different action spaces and frequencies.
Appendix B Architectures
BEAST-F is a new VLA model that integrates BEAST with the pretrained VLM Florence-2 [47]. Florence-2 is a compact vision-language model with 0.77B parameters, featuring an encoder-decoder language model architecture paired with a DaViT image encoder[59] that efficiently compresses images into sequences into just 50 tokens. Florence-2 offers two unique advantages: First, Florence-2 was pretrained with a large-scale dataset focusing on image-grounding with tasks like bounding box prediction and image segmentation. These objectives are well aligned with robotic manipulation challenges. In addition, its small size and low image token count make it computational and memory efficient and enable us to run VLA experiments on consumer-grade hardware. This combination of task-relevant pretraining and computational efficiency made Florence-2 ideal for exploring our tokenizers within practical resource constraints.
BEAST-D is a compact Transformer model that integrates discrete BEAST tokens. The architecture is shown in Figure 10.
@@ Appendix B Architectures
Appendix C Baselines Implementation
@@
Appendix D Preliminary Results with Flow Matching
We conduct an additional experiment on combining BEAST’s continuous tokens with flow-matching loss[60, 61] and a mixture-of-expert backbone[55]. The resulting model, BEAST-Flow achieves performance comparable to the Flow policy with action chunking on LIBERO-Long, and outperforms it on Calvin-ABC, while predicting only half as many tokens. The full results are presented in Figure 11. These results highlight the effectiveness of BEAST within a flow-matching framework.
LIBERO-LONGCALVIN ABC-D0505010010091.591.561.661.691.591.563.663.6Success Rate (%)FlowBEAST-Flow
Figure 11: Preliminary Results with Flow Matching.
Appendix E Hyperparameters
@@
Appendix F Real Robots Setup & Tasks
F.1 Robot System Details
@@
F.2 Tasks Description and Evaluation Metrics
@@
Appendix G Compute Resources